OpenStudy (whateveriwant):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below: f(x) = 0.69(1.03)^x Part A: Is the price of product A increasing or decreasing and by what percentage per year? please justify your answer. Part B: The table below shows the price f(t), in dollars, of product B after t years: t (number of years) 1 2 3 4 f(t) (price in dollars) 10,100 10,201 10,303.01 10,406.04

OpenStudy (whateveriwant):

Which product recorded a greater percentage change in price over the previous year? Justify your answer.

OpenStudy (whateveriwant):

please help!!!

OpenStudy (whateveriwant):

@mathmale

OpenStudy (wanell):

for B, take the price of year 2 - the price of year 1, you get 10,201 - 10,100=101 then take the price of year 3- the price of year 2, you get 10,303.01-10201=102.01 (a little bit up, right?) do the same with year 4, 3 10,406.04-10,303.01=103.03 (a little bit up again, right?)

OpenStudy (wanell):

so the price of product B increases year after year that allows me make the conclusion as above

OpenStudy (whateveriwant):

Thank you so much... can you please help me with 2 more or is that asking to much???

OpenStudy (whateveriwant):

I'm new

OpenStudy (wanell):

umm idk i guess

OpenStudy (whateveriwant):

How do I medal more than once??? :-)

OpenStudy (whateveriwant):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for three different years: Number of years 1 2 3 Option 1 (amount in dollars) 1300 1690 2197 Option 2 (amount in dollars) 1300 1600 1900 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. (2 points) Part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points) Part C: Belinda wants to invest in an option that would help to increase her investment value by the greatest amount in 20 years. Will there be any significant difference in the value of Belinda's investment after 20 years if she uses option 2 over option 1? Explain your answer, and show the investment value after 20 years for each option. (4 points)

OpenStudy (wanell):

A: linear; exponential. The change between years is constant for option 1 and not for option 2.

OpenStudy (wanell):

Yes; 4871 is significantly greater than 3000

OpenStudy (whateveriwant):

Do the links just say what the answer is?

OpenStudy (wanell):

umm no i cant really tell the the direct because mathmale gonna tell me something agine

OpenStudy (whateveriwant):

Ok, but what can I do with said numbers?

OpenStudy (whateveriwant):

@wanell

OpenStudy (wanell):

? said numbers ?

OpenStudy (whateveriwant):

in the links

OpenStudy (wanell):

ok let me go over this

OpenStudy (wanell):

Number of years 1 2 3 Option 1 (amount in dollars) 1300, 1690, 2197 Option 2 (amount in dollars) 1300, 1600, 1900) is that right

OpenStudy (wanell):

option 1 is exponential because it changes by a factor of 1.3 each year option 2 is linear because it increases by a fixed amount each year, here 300 .

OpenStudy (wanell):

part b) is to find the linear and exponential functions

OpenStudy (wanell):

for part b) Option 1: Directions: Fit data to exponential model f(n) = a * b^n if you divide 1300 by 1.3 , you will get 1000. Therefore n=0 , f(0) = 1000 so the exponential function is f(n) = 1000 * (1.3)^n Option 2: Directions: Fit data to a linear model f(n) = m*n + b if you calculate slope you get m = 300 and subtract 300 to get n=0, f(0) = 1000 So the linear model is f(n) = 300*n + 1000

OpenStudy (wanell):

option 1: y(x) = A*b^x use 2 points to find, A and b

OpenStudy (wanell):

0,1000) gives y(0) = 1000 = A*b^0 1000 = A The function is now y(x) = 1000 *b^x

OpenStudy (wanell):

(1 , 1300) second point y(1) = 1300 = 1000*b^1 1300 = 1000b b=1.3 The function is y(x) = 1000*(1.3)^x

OpenStudy (wanell):

option 2 is a line, starts at y intercept of 1000 investment, and slope of 300 dollars per year y(x) = 300x + 1000

OpenStudy (wanell):

option1 y(x) = 1000*(1.3)^x option2 y(x) = 300x + 1000 Graph them, and describe what you see.

OpenStudy (wanell):

part c, look at x=20, and go straight up, which one has a larger y value, it will be the exponential option 1.

OpenStudy (whateveriwant):

sorry for absence had to do something but please clarify @wanell @mathmale